### v1

# Local Thermal Operations and Classical Communication

Abstract
Quantum mechanics and thermodynamics are foundational to modern physics. This paper introduces a novel framework, based on local operations with classical communication, fundamental paradigm in quantum information, incorporating the preservation of thermal equilibrium: local thermal operations and classical communication (LTOCC). We introduce a hierarchy of LTOCC, ordered according to the amount of control and communication. Next, we provide a detailed analysis of symmetric local thermal operations and classical communication ($\mathcal{S}$LTOCC), including majorization-based bounds on possible transitions and properties such as generation of strong correlations from Gibbs states.
1 Introduction
Amongst the pillars on which modern physics stands, quantum mechanics and thermodynamics stand as two of the most robust ones. As sir Arthur Eddington has famously stated, “(…) if your theory is found to be against the second law of thermodynamics I can give you no hope; there is nothing for it but to collapse in deepest humiliation [1].” Similarly, despite significant controversies regarding its ontological status, quantum mechanics has withstood all experimental inquiry it has been subjected to over the century of its progress.
Recent years have seen a growth of the efforts focused on joining these two fields in what is called the quantum thermodynamics [2, 3, 4], with results applicable to the problems of quantum battery charging [5, 6, 7, 8, 9, 10, 11, 12, 13], cooling [14, 15, 16, 17, 18, 19, 20], autonomous quantum machines [21, 22, 23, 24] and more. Most importantly in the current context, however, are efforts to provide ultimate bounds on transformations of low-dimensional systems in contact with heat bath that would respect the conservation of thermal equilibrium. This approach has resulted in introducing the resource-theoretic framework of thermal operations, with wide body of related works already present within the literature [25, 26, 27, 28, 29, 30, 31, 32]
In parallel, the discovery of quantum entanglement and the related notion of quantum nonlocality lead to realization of the fundamental difference between classical and quantum distributed system, with diverse applications, such as teleportation [33], quantum key distribution [34], quantum secret sharing [35] and quantum network nonlocality [36], to name but a few. Early on it has been also realised, that local operations and classical communication (LOCC) is set of transformations, that effected on a nonlocal state do not create entanglement. Thus, a resource theory of entanglement based on LOCC as the set of free operations has been developed and studied in depth, leading to results on entanglement distillation, conversion and others.
In this work, following sir Eddington’s advice, we propose a framework for operations limited both by locality and communication restricted to classical means, as well as incorporating thermodynamic laws, thus adhering to the basic requirements of thermodynamics. For simplicity, we will refer to such operations as local thermal operations and classical communication (LTOCC). In Section 2 we recall all the notions necessary to introduce our results, including noisy and thermal operations and cooling maps, local operations and classical communication, and concepts connected to stochastic tensors and multistochasticity. In Section 3 we introduce the proper framework of LTOCC together with a hierarchy of processes requiring different amounts of communication and control. Finally, in Section 4 we focus on a detailed analysis of a restricted scheme, with additional symmetricity constraints symmetric local thermal operations and classical comminication ($\mathcal{S}$LTOCC). We discuss the problem of extremal channels, consider possible transitions in a fashion similar to the prior results for thermal operations and examine the unexpected properties of $\mathcal{S}$LTOCC such as the creation of strong correlations even for Gibbs states as inputs. Section 5 serves as a brief summary of the proposed framework and sets up a list of open problems worth further investigation.
Figure 1: Artists depiction of Alice and Bob using classical communication, in contact with local thermal baths, thus realising the backbone of LTOCC, Local Thermal Operations and Classical Communication (Courtesy of A. de Oliveira).
2 Preliminaries
2.1 Noisy operations, thermal operations and cooling maps
Let us begin by introducing the essential frameworks that describe evolution of finite-dimensional quantum states under thermodynamical constraints coming from the contact with a thermal bath characterised by inverse temperature $\beta=(k_{B}T)^{-1}$. The primary system of dimension $d$, prepared in an initial state $\rho$ is characterised by its Hamiltonian $H=\sum_{i}E_{i}\outerproduct{E_{i}}{E_{i}}$ |, where we assume the energy levels $E_{i}$ to be non-degenerate. The total system-bath initial state is assumed to be uncorrelated, $\rho\otimes\gamma_{E}$, with the bath characterised by the Hamiltonian $H_{E}$ starting in the Gibbs state defined as
$\gamma_{E}=\frac{e^{-\beta H_{E}}}{\text{Tr}(e^{-\beta H_{E}})}.$ (1)
[…]
Thermal operations framework (TOs) [25, 37, 38], with its infinite temperature version, $\beta=0$, referred to as noisy operations (NOs), works with a minimal assumption that system and bath undergo joint unitary evolution $U$ after which the bath is discarded, thus defining the channel
$\mathcal{E}(\rho)=\text{Tr}_{E}\quantity[U\quantity(\rho\otimes\gamma_{E})U^{\dagger}].$ (2)
Moreover, to ensure energy conservation, it is assumed that the unitary operation $U$ commutes with the joint Hamiltonian,
[$\commutator{U}{H\otimes\mathbbm{1}_{E}+\mathbbm{1}\otimes H_{E}}=0,$ , (3)
so it preserves the energy of the joint system.
In order to provide proper background to the main topic of this manuscript we need to consider the conditions for the convertibility of states. In the case of finite temperatures, given a pair of states $\rho$ and $\sigma$, the transition is possible if and only if there exists a thermal operation $\mathcal{E}$ such that $\mathcal{E}(\rho)=\sigma$ and $\mathcal{E}(\gamma)=\gamma$. However, as we limit ourselves to energy-incoherent states, the condition reduces to existence of a stochastic matrix $\Lambda$ acting on the respective populations $\mathbf{p}=\operatorname{diag}_{H}(\rho),\,\mathbf{q}=\operatorname{diag}_{H}(\sigma)$ such that
$\Lambda\mathbf{p}=\mathbf{q},$ $\Lambda\boldsymbol{\gamma}=\boldsymbol{\gamma}.$ (4)
In the particular case of infinite temperature, $\beta=0$, the Gibbs state reduces to the maximally mixed state, with constant populations $\eta_{i}=\frac{1}{d}$, and the Gibbs-preserving condition $\Lambda\boldsymbol{\eta}=\boldsymbol{\eta}$ defines the set of noisy operations acting on the energy-incoherent states equivalent to the set of bistochastic matrices. The state convertibility is then governed by majorization relation.
Definition 1 (Majorisation).
Given two $d$-dimensional probability distributions $\mathbf{p}$ and $\mathbf{q}$, we say that $\mathbf{p}$ majorises $\mathbf{q}$, and denote it by $\mathbf{p}\succ\mathbf{q}$, if and only if the following condition holds:
$\sum_{i=1}^{k}p_{i}^{\downarrow}\geq\sum_{i=1}^{k}q_{i}^{\downarrow}\quad\text{for all}\quad k\in\{1\dots d\},$ (5)
where $\mathbf{p}^{\downarrow}$ denotes the vector $\mathbf{p}$ rearranged in a non-increasing order.
Theorem 1 (Theorem II.1.10 of Ref. [39]).
There exists a bistochastic matrix $\Lambda$, $\Lambda\boldsymbol{\eta}=\boldsymbol{\eta}$, mapping $\mathbf{p}$ to $\mathbf{q}$ if and only if $\mathbf{p}\succ\mathbf{q}$.
The finite temperature case of $\beta>0$ has been considered in detail in [27], and here we present only the resulting relation of thermomajorisation. In order to do so, we first begin by defining slope vector $\mathbf{r}(\mathbf{p})$ with entries $r_{i}(\mathbf{p})=p_{i}/\gamma_{i}$. Then, the $\beta$-ordering of a state.
Definition 2.
Consider a $d$-dimensional population vector $\mathbf{p}$ and a corresponding Gibbs state $\boldsymbol{\gamma}$. A permutation $\pi_{\mathbf{p}}\in\mathcal{S}$ is called the $\beta$-order of the state $\mathbf{p}$ if it orders the corresponding slope vector in a non-increasing order,
$i<j\Longrightarrow r_{\pi_{\mathbf{p}}^{-1}(i)}(\mathbf{p})\geq r_{\pi_{\mathbf{p}}^{-1}(j)}(\mathbf{p}).$ (6)
Furthermore, we will call the reordered version $p^{\beta}_{i}=p_{\pi(i)}$ as the $\beta$-ordered state.
A thermomajorisation curve $f^{\,\beta}_{\mathbf{p}}:\left[0,1\right]\rightarrow\left[0,1\right]$ is defined as a piecewise linear curve composed of linear segments connecting the point $(0,0)$ and the points defined by consecutive subsums of the $\beta$-ordered form of the probability $\mathbf{p}^{\beta}$ and the Gibbs state $\mathbf{\gamma}^{\beta}$,
$\left(\sum_{i=1}^{k}\gamma^{\,\beta}_{i},\leavevmode\nobreak\ \sum_{i=1}^{k}p^{\,\beta}_{i}\right):=\left(\sum_{i=1}^{k}\gamma_{\mathbf{\pi}^{-1}_{\mathbf{p}}(i)},\leavevmode\nobreak\ \sum_{i=1}^{k}p_{\mathbf{\pi}^{-1}_{\mathbf{p}}(i)}\right),$ (7)
for $k\in\{1,\dots,d\}$.
Using the concept of thermomajorisation curves one can define the relation of thermomajorisation. Given two $d$-dimensional probability distributions $\mathbf{p}$ and $\mathbf{q}$, and a fixed inverse temperature $\beta$, we say that $\mathbf{p}$ thermomajorises $\mathbf{q}$ and denote it as $\mathbf{p}\succ_{\beta}\mathbf{q}$, if the thermomajorisation curve $f^{\,\beta}_{\mathbf{p}}$ is above $f^{\,\beta}_{\mathbf{q}}$ everywhere, i.e.,
$\mathbf{p}\succ_{\beta}\mathbf{q}\iff\forall x\in[0,1]:\leavevmode\nobreak\ f^{\,\beta}_{\mathbf{p}}(x)\geq f^{\,\beta}_{\mathbf{q}}(x)\,.$ (8)
It may happen that given two vectors, $\mathbf{p}$ and $\mathbf{q}$, are incomparable, meaning that $f^{\,\beta}_{\mathbf{p}}$ and $f^{\,\beta}_{\mathbf{q}}$ cross at some point (see Fig. 2). Furthermore, the thermomajorisation order cannot be seen as a lattice in the sense that for $\mathbf{p}$ and $\mathbf{q}$, which do not share the same $\beta$-order, there is no unique join or meet [40].
Figure 2: Example of majorization and thermomajorization curves for a pair of distributions $\mathbf{p}=(0.32,0.45,0.02,0.21),\,\mathbf{q}=(0.25,0.4,0.2,0.15)$ in $d=4$ with energy levels $E_{i}=i$. In both examples, the distribution $\mathbf{p}$ corresponding to the blue curve majorizes the distribution $\mathbf{q}$ corresponding to the orange curve, $\mathbf{p}\succ\mathbf{q},\,\mathbf{p}\succ_{\beta}\mathbf{q}$, since the blue curve lies entirely above the orange one.
We can now state the generalisation of Theorem 1 from bistochastic matrices to Gibbs-preserving matrices which provides the necessary and sufficient conditions for the existence of a thermal operation between two energy-incoherent states [27, 41].
Theorem 2 (Theorem 1.3 of Ref. [25]).
There exists a stochastic Gibbs-preserving matrix $\Lambda$, $\Lambda\mathbf{\gamma}=\mathbf{\gamma}$, mapping $\mathbf{p}$ to $\mathbf{q}$ if and only if $\mathbf{p}\succ_{\beta}\mathbf{q}$.
Bistochastic matrices correspond to thermal operations in infinite temperature. However one can also study the opposite zero temperature limit. In the A we summarize the concept of cooling maps, originating for such a study [42], in a scenario with diagonal quantum states. Furthermore in B we show an alternative derivation of cooling maps for diagonal quantum states as a zero temperature limit of Gibbs-preserving matrices.
In further discussion, following the study of thermal operations, we restrict ourselves to the states that are diagonal in the Hamiltonian eigenbasis. This lets us identify quantum states with their populations $\rho=\sum_{i}p_{i}\;|E_{i}\rangle\langle E_{i}|\Longleftrightarrow\bf{p}$.
2.2 Local operations and classical communication
Whenever considering multipartite quantum system composed of spatially separated subsystems, each with local Hilbert space $\mathcal{H}_{i}$, it is natural to limit the operations that can be effected on the total system. Such scenario is often described as "distant laboratories" paradigm. The most rigorous restriction comes from allowing the involved parties to perform operations only locally. Thus, for a bipartite system $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ the allowed operations would be limited to channels of the form $\mathcal{E}_{A}\otimes\mathcal{E}_{B}$.
It is however natural to note that distant laboratories do not necessarily mean disconnected laboratories, which leads to inclusion of classical communication, such as shared randomness or exchange of results of experiments. Together with local operations, possibly conditioned on the results of the local measurements, is referred to as local operations and classical communication (LOCC) [43]. Some of the most prominent protocols in quantum information, including quantum teleportation [44] and superdense coding [45], rely on LOCC operations; furthermore, the study of convertibility of entangled states under LOCC, which are known not to increase entanglement, is one of the cornerstones of the quantum resource theories [46, 47].
Despite their widespread use and intuitive understanding, the structure of the set of LOCC operations is intricate – we refer the reader interested in detailed description to [43], where LOCC is introduced with full rigorousness within the framework of quantum instruments. In this work, however, the most important aspect of LOCC is possibility of performing local operations conditioned on results of measurement in a distant laboratory.
2.3 Multistochastic tensors and multi stochastic Birkhoff polytope
Another important foundation for our work are stochastic tensors. Similarly, as stochastic matrix $\Lambda_{ij}$ encompasses the mapping of probability distributions, the stochastic tensor $T_{ijk}$ describe the mapping of two distributions into one,
$T(\mathbf{p},\mathbf{q})_{i}=\sum_{jk}T_{ijk}p_{j}q_{k}\leavevmode\nobreak\ ,$ (9)
and the concept can be extended for more distributions by considering tensors of higher rank. To ensure, that the operation is well-defined, all elements of the stochastic tensor must be positive, $T_{ijk}\geq 0$, and sums of its hypercolumns must be equal one, $\sum_{i}T_{ijk}=1$, for any $j,k$. Typically, physically interesting phenomena are described using restricted scenarios such as bistochasticity, which appears naturally in noisy operations or as transition probabilities arising from unitary evolution of quantum systems [48, 49] . In the case of tensors instead of one additional condition for normalisation of rows (hypercolumns in relation to a selected index), one can demand similar conditions to be satisfied by all hypercolumns, which gives rise to multistochastic tensors.
Definition 3.
A tensor $T$ is called tristochastic if $T_{ijk}\geq 0$ and $\sum_{i}T_{ijk}=\sum_{j}T_{ijk}=\sum_{k}T_{ijk}=1$ for any $i,j,k$.
The additional conditions correspond to the preservation of flat distribution $\boldsymbol{\eta}$, if it appears as any input:
$\forall\;\mathbf{p}\leavevmode\nobreak\ \leavevmode\nobreak\ T(\mathbf{p},\boldsymbol{\eta})=T(\boldsymbol{\eta},\mathbf{p})=\boldsymbol{\eta}$ (10)
The resemblance between bistochastic matrices and multistochastic tensors suggests, that many of the properties of the former should translate to the latter. However, such a connection is not entirely straightforward and becomes even more troublesome when considering quantum counterpart for both of them [50].
For example, sets of both objects are convex, forming the so-called Birkhoff polytopes [51]. All extremal bistochastic matrices, vertices of Birkhoff polytope, are permutation matrices [52] – all its entries are either zero or one. Thus one would expect, that permutation tensors, in which all entries are equal to either zero or one, also fully describe the extremal tensors. However, as proven in [53], there are also other extremal multistochastic tensors. Note, that while layers of permutation tensors are extremal bistochastic matrices, it no longer holds for other extremal multistochastic tensors.
An example of a tristochastic permutation tensor for dimension $d=3$ is provided below [54]:
$T=(T_{0,j,k},T_{1,j,k},T_{2,j,k})=\left(\begin{matrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{matrix}\right.\left|\begin{matrix}0&1&0\\ 0&0&1\\ 1&0&0\\ \end{matrix}\right.\left|\begin{matrix}0&0&1\\ 1&0&0\\ 0&1&0\\ \end{matrix}\right)\,,$ (11)
where the reader should imagine the square sub-matrices arranged in a $3\times 3\times 3$ cube, while the example of other extremal tensors is presented below
$T=(T_{0,j,k},T_{1,j,k},T_{2,j,k})=\left(\begin{matrix}0&\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&\frac{1}{2}&0\\ \frac{1}{2}&0&\frac{1}{2}\\ \end{matrix}\right.\left|\begin{matrix}\frac{1}{2}&\frac{1}{2}&0\\ 0&\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&0&\frac{1}{2}\\ \end{matrix}\right.\left|\begin{matrix}\frac{1}{2}&0&\frac{1}{2}\\ \frac{1}{2}&0&\frac{1}{2}\\ 0&1&0\\ \end{matrix}\right)\,.$ (12)
3 Framework of Local Thermal Operations and Classical Communication
Having presented all the necessary foundational ideas and concepts, we may proceed to introduction of the central idea of this work, combining spatially separated parties with thermal restrictions on the operations that they can effect – Local Thermal Operations and Classical Communication (LTOCC).
3.1 Local thermal operations and preservation of product Gibbs state
Let us consider a scenario in which Alice and Bob have two separate physical subsystems and each one of them is able to perform thermal operations. Since two subsystems are physically separated, we assume that the interactions between them are negligible. Otherwise, the parties might leverage those interactions, effectively treating two subsystems as one joint system.
Therefore the total Hamiltonian is a sum of local Hamiltonians $H=H_{A}+H_{B}$. Hence the thermal state of the entire system is a product of thermal states on each subsystem ${\gamma}={\gamma}_{A}\otimes{\gamma}_{B}$. This assumption, natural for separated subsystems, results in a crucial property: The product of any local thermal operations is thermal as well.
3.2 Classical communication and Landauer cost of measurement
The final ingredient one requires before going into the framework of LTOCC is the problem of local measurements and classical communication between the parties, as there are certain subtleties that need to be taken into account, especially the energetic cost of measuring the system in the energy eigenbasis.
The energy-incoherent quantum states $\rho=\sum_{i}p_{i}\;|E_{i}\rangle\langle E_{i}|$ can be interpreted as a probabilistic combination or a statistical ensemble of pure Hamintonian eigenstates. The measurement collapses it to a random eigenstate, thus one round of measurement is not thermal – it can project a Gibbs state onto a high energy eigenstate, resulting in an energetic gain. However, statistically, the measurements leave diagonal states invariant. Thus while performing measurement on consecutive states form a Gibbs ensemble, with respect to energy eigenbasis, $N$ times the disturbance effect of measurements vanishes on average as we go into the limit of very large number of rounds, $N\rightarrow\infty$.
The second element is the transfer and storage of information. The processing of information, according to the Landauer principle [55, 56], imposes a fundamental thermodynamic cost. Fortunately, this cost is associated with the erasure of information, which would happen between rounds or runs of protocol, on classical computers used for communication. Thus assuming that classical computers are well separated from the quantum system (except at the moment of measurement), one can disregard the contribution of Landauer cost. However, we may also limit ourselves to a finite sample and state, that after $N$ measurements the cost of erasing the memory can be estimated as $Nk_{B}T\log d$, where $d$ is a dimension of the quantum system, incurred on the side of the classical memory. After such erasure states from consecutive rounds are indistinguishable constituting the Gibbs ensemble. Thus statistically it is indeed preserved.
Finally, because statistically neither thermal operations nor measurement in energy eigenbasis can create any coherences, description by probability distributions is complete, and instead of quantum channels, we can continue our discussion using stochastic maps.
3.3 Framework outline
First, let us state the minimal requirements for any action that Alice and Bob could perform while preserving respective local equilibria. Let $M_{ij,kl}$ be a stochastic map acting on two systems [^1] In what follows, unless stated otherwise, all stochastic tensors and matrices depend, implicitly, on the inverse temperature $\beta$.. Since neither Bob nor Alice can do anything to change the Gibbs state on either of the subsystems we have:
$\forall_{k}:\sum_{i}M_{ij,kl}=:B_{j,l}^{(1,k)};\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{l}:\sum_{j}M_{ij,kl}=:B_{i,k}^{(2,l)}\leavevmode\nobreak\ \leavevmode\nobreak$ (13)
where $B^{(1,k)}$, $B^{(2,l)}$ are thermal operations for all values of $k$ and $l$ respectively. In the above equation $k,l$ are input and $i,j$ are output indexes corresponding respectively to Alice’s and Bob’s subsystems. This can also be stated in terms of input and output states as
$\forall_{k}:\sum_{i,l}M_{ij,kl}\gamma_{l}=\gamma_{j};\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{l}:\sum_{j,k}M_{ij,kl}\gamma_{k}=\gamma_{i}\leavevmode\nobreak\ .$ (14)
We will call any channel, that satisfies those two properties locally Thermal Operation ($\ell$TO).
Now we proceed with the discussion of the actual scheme, concerned with what Alice and Bob can perform using local thermal operations, measurements and communication. After each of them receives their own part of the state, Alice can perform a measurement on her subsystem and send the information to Bob who, depending on the received information, can perform different thermal operations. If we stack Bob operations conveniently in one tensor $T_{jkl}$, the discussed example corresponds to the stochastic matrix of the form
$M_{ij,kl}=\delta_{i,k}T_{jkl}\leavevmode\nobreak\ ,$ (15)
since the ensemble of Alice’s states remains unchanged. In parallel to Alice, Bob could also perform a measurement of his input state, and send the information to Alice, who can also perform stochastic operations depending on the received information. Because measurement in the energy eigenbasis does not alter the ensembles of received states we can describe this joint operation as:
$M_{ij,kl}=T_{ikl}^{(1)}T_{jkl}^{(2)}$ (16)
which we will call one round (without feedback) Local Thermal Operations and Classical communication (LTOCC). The minimal requirements, applied on (16) results in
$\forall_{l}\sum_{k}T_{ikl}^{(1)}{\gamma_{E}}_{k}={\gamma_{E}}_{i}\;,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{k}\sum_{l}T_{jkl}^{(2)}{\gamma_{E}}_{l}={\gamma_{E}}_{j}\;.$ (17)
We name any stochastic tensor satisfying one of the above requirements a thermal tensor.
This framework can be easily extended to multiple rounds without memory
$M=\prod_{a}M^{(a)}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ \leavevmode\nobreak\ M_{ij,kl}^{(a)}=T_{ikl}^{(1,a)}T_{jkl}^{(2,a)}$ (18)
or with memory, which would correspond to the growing number of indices for tensors in consecutive rounds, to describe dependencies on past measurement results. Thus replacing indices $k,l$ with multi-indices $\mathbf{k}=\quantity(k_{1},\ldots,k_{n}),\,\mathbf{l}=\quantity(l_{1},\ldots,l_{n})$ in the $n$-th round.
Another possible extension is the introduction of shared randomness (LTOCC+R). Alice and Bob can share a random variable $\lambda_{\alpha}$ and perform different LTOCC depending on its value:
$M_{ij,kl}=\sum_{\alpha}\lambda_{\alpha}T_{ikl}^{(1,\alpha)}T_{jkl}^{(2,\alpha)}$ (19)
Finally, we can impose an additional symmetry requirement. namely that if Alice received a thermal state as an input, then the result of Bob’s action would also be thermal and vice versa
$\forall_{k}\sum_{l}T_{ikl}^{(1)}{\gamma_{E}}_{l}={\gamma_{E}}_{i}\;,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{l}\sum_{k}T_{ikl}^{(1)}{\gamma_{E}}_{k}={\gamma_{E}}_{j}\;,$ (20)
and analogously for $T_{jkl}^{(2)}$.
We call any stochastic tensor satisfying both properties (20) a bi-thermal tensor. The above scheme constructed from a bi-thermal tensor gives one round of Symmetric Thermal Operations and Classical Communication ($\mathcal{S}$LTOCC).
In this work, we focus on the building block of this framework, the (bi-)thermal tensor $A$, discussing its properties, the structure of the set of such tensors and the possibility of obtaining outcomes for given input distributions.
[…]
Figure 3: Diagram of inclusions of all above-discussed classes of quantum protocols. Each frame corresponds to a different number of rounds in the protocol.
4 Detailed results on $\mathcal{S}$LTOCC
In this section, we shift our attention to the building block of the presented framework – thermal tensor. Particularly we focus on bi-thermal tensors, with the additional requirement of symmetry, to present its wide range of properties even in such restricted case. We start the discussion by considering the geometry of bi-thermal tensors’ set to unveil the extremal operations.
Similarly to the tri-stochastic tensors, the bi-thermal tensors are defined by a finite set of linear (in)equalities (20), and thus they form a convex polytope. Moreover, to fully describe this set it is sufficient to consider only a discrete set of extremal points. Let us start with mentioning the limit of infinite temperature, $\beta=0$ – tristochastic tensors. In this case, the problem of characterizing vertices was already extensively studied in [57], where they have been divided into the permutation and non-permutation extreme points.
In the finite temperature case, $\beta>0$, the situation is more convoluted, even when restricted to standard thermal operations [29]. As presented in [58], the structure of vertices for the thermal equivalent of Birkhoff polytope depends on relations between systems energies and temperature. Thus we start the discussion by presenting a simple, yet comprehensive, example of a bi-thermal operation on a 2-level system.
Note that for two level system the defining relations of bi-thermal operations (20), determine all tensor values as a function of only one. Thus geometrically it is equivalent to a line segment. To describe its extremal points let us denote the energies by $E_{0}<E_{1}$, moreover to simplify the notation we set $g_{1,0}=e^{-\beta(E_{1}-E_{0})}$. The extremal bi-thermal tensors are:
$T_{ijk}^{(1)}=(T_{0,j,k}^{(1)}|T_{1,j,k}^{(1)})=\left(\begin{matrix}1-g_{1,0}&1\\ 1&0\\ \end{matrix}\right.\left|\begin{matrix}g_{1,0}&0\\ 0&1\\ \end{matrix}\right)\leavevmode\nobreak\ ,$ (21)
$T_{ijk}^{(2)}=(T_{0,j,k}^{(2)}|T_{1,j,k}^{(2)})=\left(\begin{matrix}1-g_{1,0}(1-g_{1,0})&1-g_{1,0}\\ 1-g_{1,0}&1\\ \end{matrix}\right.\left|\begin{matrix}g_{1,0}(1-g_{1,0})&g_{1,0}\\ g_{1,0}&0\\ \end{matrix}\right)\leavevmode\nobreak\ .$ (22)
Note that while the layers of $T^{(1)}$ are extremal thermal operations, it no longer holds for the layers of $T^{(2)}$. This is because the tensor $T^{(2)}$ can be realized as $T^{(1)}$ followed by a thermal swap $S^{(\beta)}$, or as a $T^{(1)}$ pre-processed by thermal swap on one of subsystems:
$T_{ijk}^{(2)}=\sum_{l}S_{il}^{(\beta)}T_{ljk}^{(1)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ T_{ijk}^{(2)}=\sum_{l}T_{ilk}^{(1)}S_{lj}^{(\beta)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ T_{ijk}^{(2)}=\sum_{l}T_{ljl}^{(1)}S_{lk}^{(\beta)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where }\leavevmode\nobreak\ S^{(\beta)}=\left(\begin{matrix}1-g_{1,0}&1\\ g_{1,0}&0\end{matrix}\right)\leavevmode\nobreak\ .$ (23)
This interconnection between extremal tensors is standard for tri-stochastic tensors [57], however in finite temperature it results in one tensor being "more-thermalized" than the second one and multiple aprication of thermal swaps generally leads to non-extremal tensors.
Figure 4: Attainable states $\mathbf{r}=T(\mathbf{p},\mathbf{q})$ in dimension $d=2$ using LTOCC lie between the extreme surfaces generated by $T_{1}$ (red-orange) and $T_{2}$ (blue-cyan). Each distribution is characterized by its first coefficient, eg. $\mathbf{p}=(p,1-p)$, and the extremal values are denoted by shaded lines. Although fully symmetric for $\beta=0$, the symmetry is lost for finite temperatures, $\beta>0$.
4.1 Approach to extremal points for high temperatures
To tackle the general structure of bi-thermal operations’ set in dimensions $d>2$ we start by considering the scenario of large temperatures as a deviation from $\beta=0$. To study the form of extremal bi-thermal tensors one can consider its series expansion around tri-stochastic tensors:
$T_{ijk}=\widetilde{T}_{ijk}+\beta A_{ijk}^{(1)}+O(\beta^{2})\leavevmode\nobreak\ .$ (24)
In the limit of large temperature, we may expand the thermal state as
${\gamma_{E}}_{i}=\frac{e^{-\beta E_{i}}}{\sum_{j}e^{-\beta E_{j}}}=\frac{1}{d}\left(1+\beta\frac{\sum_{m}E_{m}}{d}\right)-\beta\frac{E_{i}}{d}+O(\beta^{2})$ (25)
Thus in the first order, the conditions for bi-thermality (20) lead to:
$\sum_{i}A_{ijk}^{(1)}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{j}A_{ijk}^{(1)}=\sum_{j}\tilde{T}_{ijk}E_{j}-E_{i},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{k}A_{ijk}^{(1)}=\sum_{k}\tilde{T}_{ijk}E_{k}-E_{i}$ (26)
Moreover, we must ensure the bi-thermal tensor $T_{ijk}$ is non-negative. Thus treating $A_{ijk}^{(1)}$ as tangent vector in the point $T_{ijk}$, this requirement corresponds to:
$A_{ijk}^{(1)}\geq 0\leavevmode\nobreak\ \text{ if }\leavevmode\nobreak\ \tilde{T}_{ijk}=0.$ (27)
The condition (27) does not restrict the values of all tangent vector $A_{ijk}^{(1)}$ elements since not all directions from $\tilde{T}_{ijk}$ are prohibited. In particular $A_{ijk}^{(1)}$ can always point in the centre of set of bi-thermal tensor with arbitrary amplitude. However, we are interested in the extremal bi-thermal tensors. Thus we focus on extremal tensor $T_{ijk}$ and extremal tangent vectors $T_{ijk}^{(1)}$, which saturate as many of (27) inequalities as possible.
This approach not only lets us recover the first order of extremal tensors expansion in dimension $d=2$ (21),(22), but also find the first order expansion in dimension $d=3$ around an exemplar permutation tensor (11), which yields $67$. The solutions, together with the code used to obtain them are presented in the github repository [59].
We note the enormous growth of extremal tensors’ number in finite temperature, even in such a small local dimension. This indicates that the complexity of the set of bi-thermal operations is immense even in comparison with standard thermal operations, for which in local dimension $d=3$ there were no more than $3$ extremal operations corresponding to one permutation [58].
4.2 Zero temperature
The opposite extremal case of zero temperature $\beta\to\infty$ turns out to be surprisingly simple. As stated in [42], for temperatures low enough, if the thermal bath has any spectral gap, effectively its entire population lies in the ground state. In such a scenario, thermal operations reduce to cooling maps defined by condition $\Lambda_{ij}=0$ if $i>j$. Moreover, in the appendix B, we present an alternative derivation of cooling maps, as a zero temperature limit $\beta\to\infty$ of thermal operations.
Knowing the limit form of thermal operations in zero temperature, we can combine them into bi-thermal operations.
Theorem 3.
In zero temperature, assuming a non-degenerated Hamiltoninan spectrum, the set of bi-thermal operations simplifies to a polytope of stochastic matrices satisfying $T_{ijk}=0$ if $i>j$ or $i>k$. The extremal operations correspond to tensors with 00 and $1$ entries, restricted only by the abovestated condition.
Hereafter we will call the zero temperature version of bi-thermal operations the bi-cooling operations.
Proof.
From the discussion of thermal operations in the zero temperature $\beta\to\infty$, we deduce, that the bi-thermality conditions (20) changes to
$T_{ijk}=0\text{ if }i>j\text{ or }i>k\leavevmode\nobreak\ .$ (28)
Thus effectively the columns of the tensor $T_{ijk}$ are decoupled and can be treated as independent probability vectors subject only to the condition (28). From this, the form of extremal tensors, as a combination of extremal columns, follows immediately. ∎
We note, that the set defined by the limiting form of bi-thermality conditions (28), might be larger than the limit of the set of bi-thermal tensors. However, we choose the first one as bi-thermal tensors in zero temperature, since in the discussed framework the fundamental physical actions are thermal operations. In the limit of zero temperature, one can combine them into any tensors satisfying conditions (28).
4.3 Explicit construction of a family of extreme bi-thermal tensors in low temperatures.
Although the general construction of extremal bi-thermal tensor seems unattainable, based on the results introduced in [58] we may construct a family of extreme bithermal tensors, which exist for every Hamiltonian when the temperature is low enough. Hereafter, for the sake of clarity, we adopt the notation $g_{i,j}=e^{-\beta(E_{i}-E_{j})}$ [58]. In dimension $2$ our example is just an extremal tensor (21)
$T_{ijk}|_{d=2}=(T_{i,0,k}|T_{i,1,k})_{d=2}=\left(\begin{matrix}1-g_{1,0}&1\\ g_{1,0}&0\\ \end{matrix}\right.\left|\begin{matrix}1&0\\ 0&1\\ \end{matrix}\right)\leavevmode\nobreak\ ,$ (29)
Note, that we sliced the tensor differently so that each slice is now a thermal operation that Bob may perform. In dimension $3$ the structure is given as follows
$T_{ijk}|_{d=3}$ $=(T_{i,0,k}|T_{i,1,k}|T_{i,2,k})|_{d=3}=\left(\begin{matrix}1-g_{1,0}-g_{2,0}&1&1\\ g_{1,0}&0&0\\ g_{2,0}&0&0\\ \end{matrix}\right.\left|\begin{matrix}1&0&0\\ 0&1-g_{2,1}&1\\ 0&g_{2,1}&0\\ \end{matrix}\right.\left|\begin{matrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{matrix}\right)=$ (30)
$=\left(\begin{matrix}1-g_{1,0}-g_{2,0}&\vline&\begin{matrix}1&1\\ \end{matrix}\\ \hline\cr\begin{matrix}g_{1,0}\\ g_{2,0}\\ \end{matrix}&\vline&\mbox{\Large 0}\end{matrix}\right.\leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ 0\\ \end{matrix}&\vline&T_{i,0,k}|_{d=2}\end{matrix}\right.\leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ 0\\ \end{matrix}&\vline&T_{i,1,k}|_{d=2}\end{matrix}\right)$
And in the higher dimensions we follow the same inductive procedure
$T_{ijk}|_{d=n}=(T_{i,0,k}|\cdots|T_{i,n,k})|_{d=n}=$ (31)
$=\left(\begin{matrix}1-\sum_{i=1}^{n}g_{i,0}&\vline&\begin{matrix}1&\cdots&1\\ \end{matrix}\\ \hline\cr\begin{matrix}g_{1,0}\\ \vdots\\ g_{d,0}\\ \end{matrix}&\vline&\mbox{\Large 0}\end{matrix}\right.\leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&\cdots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\\ \end{matrix}&\vline&T_{i,0,k}|_{d=n-1}\end{matrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \right|\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdots\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&\cdots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\\ \end{matrix}&\vline&T_{i,n-1,k}|_{d=n-1}\end{matrix}\right)\leavevmode\nobreak\ .$
To guarantee the non-negativity of each element in the tensor we must demand
$\sum_{i=1}^{n}g_{i,0}\leq 1\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i=2}^{n}g_{i,1}\leq 1\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \cdots,\leavevmode\nobreak\ g_{n,n-2}+g_{n-1,n}\leq 1\leavevmode\nobreak\ .$ (32)
Each of those equations introduces some upper bound for the temperature, thus it is sufficient to satisfy the tightest of them. The above bi-thermal tensor is symmetric in the sense $T_{ijk}=T_{ikj}$, see fig. 5, thus the opposite way of slicing it would give the same formulas.
Note that each layer of this tensor is an extreme thermal operation. It has a block structure consisting of identity and an extremal thermal operation presented in [58] (see fig. 5). Since none of the tensor layers can be decomposed as a convex combination of thermal operations, the entire tensor itself cannot be either, thus it is extremal.
Furthermore, each layer cannot be decomposed as a product of two other thermal operations, as demonstrated in [58], so this tensor cannot be obtained from anything "simpler" using local thermal pre- and post-processing. Finally in the limit of zero temperature, the bi-thermal tensor (31) simplifies to a bi-cooling tensor of least possible cooling i.e.
$T_{ijk}|_{d=n,T=0}=\left\{\begin{matrix}1&\text{ if }i=\min(j,k),\\ 0&\text{ otherwise},\end{matrix}\right.$ (33)
thus in each column, the populations are placed at the highest possible temperature.
Figure 5: Graphical representation of extremal bi-thermal tensor (31) in dimension $d=4$ as a cube and its separation into consecutive slices. Grey cells correspond to value $1$, red-to-blue to other other nonzero values, and empty cells to 00.
4.4 Local action
Having discussed the general structure of the set of bithermal operations, including methods for obtaining the extreme points, we shift our attention to the images of pairs of states under $\mathcal{S}$LTOCC. We start our discussion by considering local results. In other words, we will consider what Alice can obtain in her subsystem, excluding the created correlations with Bob’s subsystem. Thus, we will focus on one (bi-)thermal tensor, rather than the entire scheme. We stress that although the entire sub-section is presented from the perspective of (bi-)thermal tensors, the presented result holds also for tristochastic tensors.
It turns out that there is a strong connection between the images of (bi-)thermal tensors and thermomajorization. In the following theorem in square brackets, we present its extended version for bi-thermal tensors.
Theorem 4.
Let $T$ be a [bi-]thermal tensor, $\forall_{q}$ $T({\mathbf{\gamma}},\mathbf{q})={\mathbf{\gamma}}$ [and $T({\mathbf{q},\mathbf{\gamma}})={\mathbf{\gamma}}$], if and only if for any input distrstibutions $p,q$ the output is a probability distribution satisfying thermomajorization relations $\mathbf{p}\succ_{(\beta)}T(\mathbf{p},\mathbf{q})$ [and $\mathbf{q}\succ_{(\beta)}T(\mathbf{p},\mathbf{q})$].
Proof.
Let $T$ be a thermal tensor then for any distribution $\mathbf{q}$, $T(\cdot\leavevmode\nobreak\ ,\mathbf{q})$ is a thermal operation, thus the implication in one direction follows.
To prove the implication in the opposite direction, we note that any layer of a tensor $B^{(j)}=T(\cdot,e_{j})$ is a matrix such that $\mathbf{p}\succ_{(\beta)}B^{(j)}(\mathbf{p})$. Thus each layer $B^{(j)}$ preserves a Gibbs state. Moreover, its non-negativity and normalization can be proven analogically as in [60] (theorem A.4). Since each layer $B^{(j)}$ is a thermal operation, $T$ must be a thermal tensor.
In the case of bi-thermal tensors, one only needs to repeat the above arguments replacing the first with the second input state. ∎
As one can see, the problem of majorization between inputs and outputs is very similar to the scenario with thermal operations. On the other hand, the problem of finding a tensor, that may transform a pair of distributions into a desired one is much more complicated. The majorization of possible output by inputs is only the necessary conditions for possible mapping.
Lemma 1.
Let $\mathbf{p},\mathbf{q},\mathbf{r}$ be three probability distributions, then there exists a bi-thermal tensor $T$ mapping $\mathbf{p},\mathbf{q}$ into $\mathbf{r}$: $\mathbf{r}=T(\mathbf{p},\mathbf{q})$ if and only if $\mathbf{r}$ is majorized by a convex combination of majorization curves of $T^{(i)}(\mathbf{p},\mathbf{q})$ for all extremal tensors $T^{(i)}$.
Proof.
Suppose some combination of majorization curves for $T^{(i)}(\mathbf{p},\mathbf{q})$ majorizes $\mathbf{r}$. In that case, one can perform a convex combination with the same weights between appropriate extremal tensors $T^{(i)}$, obtaining tensor $\tilde{T}$ such that $\tilde{T}(\mathbf{p},\mathbf{q})\succ_{(\beta)}\mathbf{r}$. Thus additional thermal post-processing is sufficient to construct the desired tensor. Note that such post-processing can be included in bi-thermal tensor without changing its defining properties.
On the other hand, if there is no convex combination of majorization curves that majorizes $\mathbf{r}$, there is also no combination of extremal tensors which would lead $\mathbf{p},\mathbf{q}$ into state distribution majorizing $\mathbf{r}$ thus there is no desired tensor. ∎
From the above lemma, we can deduce the following observations, which simplify the discrimination of possible transforming bi-thermal tensors.
Observation 1.
Let $l$ be a curve obtained by, at each point, taking a maximum over majorization curves for $T^{(i)}(\mathbf{p},\mathbf{q})$, with $T^{(i)}$ being once again an extremal tensors. If $l$ does not majorize the curve corresponding to $\mathbf{r}$, then there exists no tristochastic (bi-thermal stochastic) tensor $T$, such that $T(\mathbf{p},\mathbf{q})=\mathbf{r}$.
Observation 2.
(Numerical) For tristochastic tensors, it is sufficient to consider only extremal permutation tensors.
The heuristic argument in favour of Observation 2 is as follows. The only tensors that map all pairs of basis vectors $e_{i},e_{j}$ into basis vectors are permutation tensors. Thus for any pair of input basis vectors its image under any permutation tensor majorizes its image under any tristochastic tensor. Since all finite probability distributions are convex combinations of basis vectors, we expect that the dominance of permutation tensors holds in the general case.
Fortunately, there exists also a stronger necessary condition for the existence of bi-thermal tensor $T(\mathbf{p},\mathbf{q})\ =\mathbf{r}$ for given distributions $\mathbf{p},\mathbf{q},\mathbf{r}$. To present it let us introduce the following notion.
For any distribution $p$ let us define its projection $\tilde{\mathbf{p}}$ onto the boundary of the probability simplex, which we will refer to as boundary distribution of $\mathbf{p}$,
$\tilde{\mathbf{p}}=\frac{1}{1-w_{p}}\mathbf{p}-\frac{w_{p}}{1-w_{p}}{\gamma},$ (34)
with $w_{p}=\min_{i}\frac{p_{i}}{{\gamma}_{i}}$; note that $\tilde{p}$ has at least one zero entry.
Theorem 5.
Let $\mathbf{p},\mathbf{q},\mathbf{r}$ be three probability distributions, and $\tilde{\mathbf{p}}$, $\tilde{\mathbf{q}}$ boundary distributions of $\mathbf{p},\mathbf{q}$. Moreover, let us define
$\overline{\mathbf{r}}=\frac{1}{(1-w_{p})(1-w_{q})}\mathbf{r}-\frac{1-(1-w_{p})(1-w_{q})}{(1-w_{p})(1-w_{q})}\gamma$ (35)
Then the necessary condition for $\mathbf{r}$ to be achivable by a bi-thermal tensor form $\mathbf{p},\mathbf{q}$ is not only $\mathbf{p}\succ_{\beta}\mathbf{r}$ and $\mathbf{q}\succ_{\beta}\mathbf{r}$ but also $\tilde{\mathbf{p}}\succ_{\beta}\overline{\mathbf{r}}$ and $\tilde{\mathbf{q}}\succ_{\beta}\overline{\mathbf{r}}$.
Proof.
Assume that there exists some bi-thermal tensor such that $T(\mathbf{p},\mathbf{q})=\mathbf{r}$. Then by the definition of boundary distributions:
$T(\mathbf{p},\mathbf{q})$ $=T\left((1-w_{p})\tilde{\mathbf{p}}+w_{p}{\gamma},(1-w_{q})\tilde{\mathbf{q}}+w_{q}{\gamma}\right)$ (36)
$=(1-w_{p})(1-w_{q})T(\tilde{\mathbf{p}},\tilde{\mathbf{q}})+\left[1-(1-w_{p})(1-w_{q})\right]{\gamma}$
Therefore the image of $\tilde{\mathbf{p}},\tilde{\mathbf{q}}$ is exactly $\overline{\mathbf{r}}$, hence it must be a probability distribution and $\tilde{\mathbf{p}}\succ_{\beta}\overline{\mathbf{r}}$, $\tilde{\mathbf{q}}\succ_{\beta}\overline{\mathbf{r}}$. ∎
Thus we effectively boosted the requirements proceeding from thermomajorization relations, leveraging the fact, that the allowed operations are bi-thermal.
Figure 6: The allowed region $\mathbf{r}=T(\mathbf{p},\mathbf{q})$ in local dimention $d=3$. The two input states are denoted by black dots. All images of $T(\mathbf{p},\mathbf{q})$ are convex combinations of extremal images, thus they lie inside the red border. The green region is the upper bound on states accessible via bi-thermal operation according to Theorem 5. (a) In infinite temperature $\beta=0$ images via permutation tensors and other extremal tensors are represented by red and blue dots, respectively. The gray dots correspond to the images of $p$ and $q$ under permutation pre-processing which does not affect the allowed region. (b) In zero temperature, $\beta\rightarrow\infty$ images via extremal tensors are denoted in red.
4.5 Action on separable states
Finally, we return to the scheme of one round of $\mathcal{S}$LTOCC between two parties to demonstrate its promising properties. Hereafter we restrict ourselves to a product input state.
First and foremost we stress that $\mathcal{S}$LTOCC operation $M_{ij,kl}=T_{ijk}^{(1)}T_{jkl}^{(2)}$ is not a thermal operation. It can induce correlations even on Gibbs ensemble $\gamma=\gamma^{(1)}\otimes\gamma^{(2)}$, which can be most clearly seen in the case of extremal operations in large or small temperatures.
Lemma 2.
Consider the limit of infinite $\beta\to 0$ or zero $\beta\to\infty$ temperature, and $\mathcal{S}LTOCC$ operation consisting of two copies of the same extremal bi-thermal channel $M_{ij,kl}=T_{ijk}T_{jkl}$, which entries are either 00 or $1$. Then for any product input state $\mathbf{p}\otimes\mathbf{q}$ the output state $\mathbf{\pi}=M(\mathbf{p}\otimes\mathbf{q})$ is maximally correlated.
Proof.
Let us denote $T(\mathbf{p},\mathbf{q})=\mathbf{r}$, then, since in each column ($\forall_{j,k}$) of $T_{ijk}$ there is only a single entry equal to $1$ and the rest of them are 00, we get
$\pi_{ij}=\sum_{kl}M_{ij,lk}p_{k}q_{l}=\sum_{kl}T_{ikl}T_{jkl}p_{k}q_{l}=\sum_{kl}\delta_{i,j}A_{ijk}p_{k}q_{l}=\delta_{ij}r_{i}\leavevmode\nobreak\ .$ (37)
Thus the output distribution is diagonal. Thus by measuring one output, one knows about the result of the second output as well. Formally, while calculating conditional entropy between outputs one obtains
$H(A|B)=H(B|A)=-\sum_{ij}\pi_{ij}\ln\left(\frac{\pi_{ij}}{r_{i}}\right)=-\sum_{ij}\delta_{i,j}r_{i}\ln\left(\frac{r_{i}}{r_{i}}\right)=-\sum_{i}r_{i}\ln(1)=0.$ (38)
The following result can be modified to encompass also some tensors in non-zero temperatures, for example, the extremal tensor at low temperatures given in (31). Notice that its last layers are identities $T_{i,j,0}=T_{i,0,k}=\mathbb{I}$. Therefore one round $\mathcal{S}$LTOCC constructed form it $M_{ij,kl}=T_{ikl}T_{jkl}$ acting on the states $\mathbf{p},\mathbf{q}$ one of which is maximally excited $\mathbf{q}=(0,\cdots,1)$ gives
$M(p,q)=\sum_{kl}T_{ikl}T_{jkl}p_{k}q_{l}=\sum_{k}T_{ik0}T_{jk0}p_{k}=\sum_{k}\delta_{i,k}\delta_{j,k}p_{k}=\delta_{i,j}p_{i}\leavevmode\nobreak\ ,$ (39)
which is once again a maximally correlated state. Moreover, its mutual information is equal to the entropy $S(p)$, so it is restricted only by $S(p)\leq\ln(n)$.
5 Summary and Outlook
In this work, we introduce a novel framework where parties operate under the distant laboratories paradigm, subject to thermodynamic constraints, coining the notion of local thermal operations and classical communication (LTOCC). We begin by presenting a hierarchy of classes of allowed operations, categorized by the level of control and the complexity of communication between the parties. A particular focus is placed on a specific subclass called symmetric LTOCC or $\mathcal{S}$LTOCC, restricted to two inputs. We explore its properties and demonstrate the complexity of this set of operations. We propose a method for generating extreme points at high temperatures by perturbing known extreme points in infinite temperature, and a concrete example of an extreme operation is provided, which exists at sufficiently low temperatures.
Next, we investigate methods to approximate accessible transformations using well-studied thermomajorization relations between the input and output states. These relations provide insight into how classical thermal constraints govern possible transformations. We also examine the impact of these operations on the joint state between the communicating parties, highlighting the high degree of correlation that emerges during the execution of the operation. This underscores the intricate interplay between thermal equilibrium and communication.
The results presented here lay the foundation for a new framework that combines two established concepts: the distant laboratories setting and thermodynamic limitations on operations. Given the novelty of this framework, several key questions remain open for future research. The first is a deeper exploration of non-symmetric and multi-round scenarios, which may allow for a broader range of accessible state transitions. Another important question is the introduction of coherences into the framework. This is a challenging problem, as coherent thermal operations are difficult to describe in closed form even for a single qubit [61, 40]. The distinction between thermal operations and Gibbs-preserving operations which emerge on the quantum level further complicates this analysis.
Moreover, in the face of recent developments at the frontier of quantum thermodynamics and nonlocality people have begun to investigate interplays between characteristic features of both subfields, such as possibility of generating entanglement using out-of-equilibrium states [62], detection of nonlocal properties of states using their thermodynamic properties [63] or ever-growing body of literature concerned with steady-state entanglement from autonomous machines operating in thermal framework [64, 65]. In all of the above cases, we believe that considering the problem within the LTOCC framework would further understanding of both thermodynamical and informational costs incurred throughout the implementation of the protocol.
Finally, the introduced framework could be applied to quantum communication protocols, potentially revealing new thermodynamic trade-offs, such as those between system temperature and achievable key rates or thermodynamical entropy of states and informational entropy of shared randomness. Such analysis would allow for a better understanding of thermally optimal operations in the context of protocols such as entanglement generation, cooling or work extraction. However, this would require studying measurements that are not restricted to the energy eigenbasis, raising questions about the inherent thermodynamic costs. These trade-offs between information-theoretic gains, like key rates, and thermodynamic costs, such as work required, represent another avenue for future research.
All of these questions, while significant, fall outside the scope of this initial presentation of the LTOCC framework. They should be viewed as a roadmap for future studies, aimed at expanding and refining this new theoretical approach.
Acknowledgments
The authors thank A. de Oliveira Jr., Jeongrak Son, Kamil Korzekwa and Karol Życzkowski for useful discussions and comments concerning the paper. Furthermore, we thank A. de Oliveira Jr. for kindly providing us with the artistic impression in Fig. 1.
RB acknowledges support by the National Science Centre, Poland, under the contract number 2021/03/Y/ST2/00193 within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733.
Appendix A Cooling maps
In this Appendix, we discuss the theoretical framework designed to grasp thermal operations in low temperatures: cooling maps [42]. Although the original construction was concerned with the general scenario of mixed quantum states, we let ourselves limit the discussion to diagonal states, similarly as in the rest of the paper.
Consider a quantum system together with a thermal bath with a nonzero spectral gap. If the temperature tends to zero, the contribution to the bath’s Gibbis state from excited states decreases exponentially. Thus for low enough temperatures, one may approximate
$\mathbf{\gamma}\approx|E_{0}\rangle\langle E_{0}|\leavevmode\nobreak\ .$ (40)
Therefore the energy-preserving interaction with such a bath (3) can only deexcite the system of interest. Indeed, if we consider a diagonal input state with populations $\mathbf{p}=\operatorname{diag}_{H}(\rho)$, the action of the cooling map does not create coherences and modify the populations according to [42]
$\mathbf{q}=P\,\mathbf{p}$ (41)
where $P$ is upper-triangular (UT) stochastic matrix: $P_{j,k}=0$ if $j>k$.
Similarly as for bistrochastic matrices and thermal operations, one can define a suitable version of majorization also for cooling maps. However this time the deexcitation properties of cooling maps naturally fix the order of basis states, so there are no rearrangements.
Definition 4 ([42]).
Given two $d$-dimensional probability distributions $\mathbf{p}$ and $\mathbf{q}$, we say that $\mathbf{p}$ UT-majorises $\mathbf{q}$, and denote it by $\mathbf{p}\succ_{UT}\mathbf{q}$, if and only if the following condition holds:
$\sum_{i=1}^{k}p_{i}\geq\sum_{i=1}^{k}q_{i}\quad\text{for all}\quad k\in\{1\dots d\}.$ (42)
This definition lets us state the counterpart of Theorems 1 and2 in zero temperature:
Theorem 6 (Theorem 1 of Ref. [42]).
There exists a cooling map, corresponding to upper triangular matrix $P$, $P\mathbf{\gamma}=\mathbf{\gamma}$, mapping $\mathbf{p}$ to $\mathbf{q}$ if and only if $\mathbf{p}\succ_{UT}\mathbf{q}$.
Appendix B Alternative derivation of cooling maps
In this Appendix we present an alternative derivation of cooling maps for diagonal states as a thermal operation in the limit of zero temperature $\beta\to\infty$.
Theorem 7.
In the zero temperature limit $\beta\to\infty$, assuming nondegenerated Hamiltonian spectrum, the stochastic matrix $\Lambda$ corresponds to thermal operation if and only if $\Lambda_{i,j}=0$ for $i>j$.
The extremal operations correspond to matrices with 00 and $1$ entries, restricted only by the condition $\Lambda_{ij}=0$ if $i>j$.
Proof.
We start the proof by discussing the necessity of the above condition, to later switch to its sufficiency. Let us divide the populations of the Gibbs ensemble into two sets. The first are those corresponding to energies smaller of equal $E_{m}$ and the other correspond to larger energies. Since the operation $\Gamma$ is thermal, Gibbs ensemble is an invariant state, $\Gamma{\gamma}={\gamma}$, so the flow of populations between those two sets, described by $\Gamma$ elements must balance, as shown below
$\sum_{i>m}\sum_{j\leq m}\Lambda_{ij}{\gamma}_{j}=\sum_{i>m}(\sum_{j}\Lambda_{ij}{\gamma}_{j}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})=$ (43)
$=\sum_{i}({\gamma}_{i}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})-\sum_{i\leq m}({\gamma}_{i}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})=1-\sum_{j>m}{\gamma}_{j}-\sum_{i\leq m}{\gamma}_{i}+\sum_{i\leq m}\sum_{j<m}\Lambda_{ij}{\gamma}_{j}$
$=\sum_{i\leq m}\sum_{j<m}\Lambda_{ij}{\gamma}_{j}$
* To obtain the bounds on $\Lambda_{ij}$ elements, we may extrapolate the populations of Gibbs states using two inequalities: ${\gamma}_{i}\geq{\gamma}_{m}=e^{-\beta E_{m}}$ for $i\leq m$, and ${\gamma}_{i}\leq{\gamma}_{m+1}=e^{-\beta E_{m+1}}$ for $i>m$. Thus the above equality can be changed into
$\sum_{i>m}\sum_{j\leq m}\Lambda_{ij}e^{-\beta E_{m}}$ $\leq\sum_{i\leq m}\sum_{j>m}\Lambda_{ij}e^{-\beta E_{m+1}}$ (44)
$\sum_{i>m}\sum_{j\leq m}\Lambda_{ij}$ $\leq(\sum_{i\leq m}\sum_{j>m}\Lambda_{ij})e^{-\beta(E_{m+1}-E_{m})}.$
Since the right-hand side of the equation tends to zero and all terms on the left-hand side are positive we obtain the desired result.
It turns out, that, the condition from this Lemma is also sufficient. This can be seen by considering the channel, which corresponds to a detailed balance of Gibbs state, ie $\Lambda_{ij}{\gamma}_{j}=\Lambda_{ji}{\gamma}_{i}$, without any summation involved. Each such stochastic channel is by its construction also thermal. However, in the limit of zero temperature $\beta=\infty$, the condition of detained balance $\Lambda_{ij}=\Lambda_{ji}e^{-\beta(E_{i}-E_{j})}$, is always satisfied given $\Lambda_{ij}=0$ if $i>j$. Hence the freedom of all nonvanishing $\Lambda$ elements is restricted only by stochasticity.
Because the condition $\Lambda_{ij}=0$ if $i>j$ does not couple the elements of $\Lambda_{ij}$, we can treat each column of $\Lambda$ separately as a probability vector which immediately leads to the extremal points. ∎